OpenStudy (anonymous):

when a 2kg block hangs at rest from an ideal spring extended 6 cm from its relaxed length.an upward external force is then applied to the bloc to move it upward a distance of 16 cm.while the block is moving upward,what would be the work done by the spring?

OpenStudy (anonymous):

T=2*3.1416*\[\sqrt{l/g}\]=.49=\[\approx.5\] strech,l=.06m mass,M=2 kg work,w=.5*k* y^2 spring constant, k=2*3.1416/(T*M)=2*3.1416/(.5*2)=6.28 hence w=.5*6.28*.16^2=.08 j